v=4/3*22/7*(5 3/4)^3
1
Expert's answer
2012-09-25T08:48:34-0400
5 3/4=23/4

4/3*22/7*(5 3/4)^3=4/3*22/7*(23/4)^3=(4*22*23^3)/(3*7*4^3)=(22*23^3)/(21*16)=(11*23^3)/(21*8)=72,4.